# Adapting Mouse Pathological Model to Human Glomerular Lesion Segmentation

### v1

@@ Abstract
Keywords: Segmentation, Glomerulus, Whole-slide image, Glomerular lesion
@@ 1 Introduction
Recent advancements such as  Isensee et al. (2021); Huang et al. (2017); Dolz et al. (2018) in image processing and deep learning have significantly influenced the enhancement of computer-assisted analysis of medical images , particularly in the quantification of animal tissues via WSI. Within the domain of renal pathology, which is characterized by its unique challenges in pathology image analysis, the glomerulus is identified as critical functional units essential for clinical assessments, according to  Jiang et al. (2021); Han et al. (2023); Østergaard et al. (2020).
Considerable research including Janowczyk and Madabhushi (2016); Komura and Ishikawa (2019) has investigated the use of deep learning in the quantification of glomeruli within the field of renal pathology, employing technologies such as convolutional neural networks (CNNs) like Gadermayr et al. (2019); Esteva et al. (2019); Wang et al. (2019); Kamnitsas et al. (2017). Interestingly, for glomerulus segmentation task, mouse-to-human transfer learning has been proven to be practical by Souza et al. (2023); Østergaard et al. (2020). Nonetheless, these studies like Yang et al. (2022) have devoted relatively little attention to the specific area of glomerular lesions. Glomerular lesions, which signify glomerular damage, serve as primary markers for various renal pathologies, according to Saikia et al. (2023). Nonetheless, the segmentation of glomerular lesions, owing to their minuscule size and the variability encountered in intra- and inter-observer analyses, presents a more formidable challenge than the segmentation of healthy glomeruli. The analysis of glomerular lesions in humans is a fundamental aspect of nephrology, offering critical insights for diagnosis, understanding disease mechanisms, guiding treatment decisions, assessing prognosis, preventing kidney failure, and facilitating research into new treatments. Moreover, very few, if any, publicly accessible datasets are available for training segmentation models, which include a broad spectrum of lesion types in mouse kidney pathology images.
@@
Figure 1: Overview. This figure provides an overview of transfer learning for glomerular segmentation from mouse to human, highlighting the advantages of using mouse samples compared to human samples.
@@
Figure 2: Lesion Class.This figure shows fine-grained classes of mouse and human glomerular images that are stained with PAS. Glomeruli with lesions: global sclerosis(GS), haylinosis(HN), mesangial lysis(ML), microaneurysm(MA), nodular sclerosis(NS), and segmental sclerosis(SS).
@@ 2 Method
We introduce the GLAM network, inspired from the research by Deng et al. (2022), tailored for our glomerular lesion segmentation tasks. GLAM Zero-shot network is a unified segmentation network that leverages a residual U-Net architecture to segment various lesion classes within partially labeled images for pathology analysis. The comprehensive structure of the GLAM Zero-shot network is detailed in Figure referred to as Fig.3.
Unlike from fully labeled dataset, each image included in partially labeled dataset contains the annotations of only a specific class of glomerular lesion. We assign each segmentation on each class of lesion as a task for task-awareness for task-awareness, by encoding the task as a $m$-dimensional one-hot vector by Chen et al. (2017). The $m$ is the number of lesion classes, The encoding calculation for $T_{k}$, a class-aware vector of $i$th class of lesion is shown as follows:
@@
Dynamic filter generation was introduced by Zhang et al. (2021) to generate the kernels specialized to a particular class of lesion segment tasks. The image feature $F$ is aggregated by a global average pooling (GAP) with $T_{k}$. The kernel parameters $\omega$ are computed as follows:
@@ 3.1 Data description
The datasets utilized in this study, as detailed in Table.1, encompass both a mouse glomerular lesion dataset and a human glomerular lesion dataset. The mouse dataset collect patches from 17 whole-slide images (WSIs) obtained from various mouse kidneys, spanning 6 lesion classes. The human dataset consists of patches derived from WSIs of 19 patients, also spanning 6 lesion classes. Initially, each patch in both datasets measured 1024x1024 pixels, but they were subsequently resized to 512x512 pixels. Each patch contains at least one type of lesion. Given the differences in glomerulus sizes between humans and mice, we conducted our analysis at different magnifications: 40x for human samples and 80x for mouse samples, ensuring similar visualization and assessment of the lesions.
Table 1: Data Collection
Samples Size Scale #Patch GS HN ML MA NS SS Sum Mouse $512^{2}$ 80×80$\times$80 × 346 187 73 198 328 193 1325 Human $512^{2}$ 40×40$\times$40 × 370 36 23 57 229 72 787
@@ 3.2 Implementation details
We designed the image pool by incorporating strategies from Cycle-GAN by Zhu et al. (2017) for each category, with a batch size of four and an image pool size of eight. Excess images in the pool, when exceeding the batch size, were removed and input into the network.
@@
We compared the introduced network to baseline models, including: (1) multiple individual U-Net models (U-Nets) by  Ronneberger et al. (2015), (2) multiple individual DeepLabv3 models (DeepLabv3s) by  Lutnick et al. (2019), and (3) a multi-class segmentation model for partially labeled datasets by  González et al. (2018) for renal pathology quantification. Additionally, the performance of the network was evaluated against transformer baselines (4) a hybrid neural network architecture that combines the Swin Transformer with the U-Net transformer encoder for enhanced medical image segmentation (Swin-Unetr) by Hatamizadeh et al. (2021), and (5) an approach to semantic segmentation based on the Vision Transformer (Segmenter) by Strudel et al. (2021). All of the parameter settings are followed by original paper.
Figure 4: Qualitative Results of different methods on human glomerular lesions. This figure displays the qualitative outcomes of various segmentation methods for four classes of glomerular lesions. The first column features the original, unannotated images, while the second column shows the manual segmentation results. Subsequent columns belong to three section: ”Mouse-to-human zero-shot transfer learning (Zero-shot)”,”Mouse-to-human few-shot transfer learning (Few-shot)” and ”Human-in-domain”, where VM refers to model selection utilizing validation results from mouse data, whereas VH indicates model selection based on validation results from human data. The bold mark indicates the best performance.
@@ 4 Results
Method Global sclerosis Haylinosis Mesangial lysis Microaneurysm Dice HD MSD Dice HD MSD Dice HD MSD Dice HD MSD U-Nets Zero-shot 70.6 185.0 51.6 49.9 352.4 317.0 49.2 521.6 441.8 50.9 273.5 119.7 DeepLabV3s Zero-shot 78.2 164.5 47.3 52.7 351.7 282.1 52.4 359.1 280.6 49.9 340.3 268.2 Multi-class Zero-shot 78.0 143.5 40.2 49.9 391.0 355.7 49.2 521.6 441.8 49.4 422.9 347.8 Swinunetr Zero-shot 81.5 147.6 37.3 50.1 319.0 251.5 49.3 225.1 94.2 49.5 255.6 109.1 Segmenter Zero-shot 82.5 146.6 36.6 50.4 267.8 224.7 49.2 521.6 441.8 50.7 251.6 121.2 GLAM Zero-shot(VM) 79.3 137.2 39.8 49.9 389.4 348.6 57.8 244.8 118.5 50.5 240.9 122.2 GLAM Zero-shot(VH) 82.0 125.3 35.9 49.9 387.5 351.9 56.2 248.9 96.8 50.2 243.8 111.4
Method Nodular sclerosis Segmental sclerosis Average Dice HD MSD Dice HD MSD Dice HD MSD U-Nets Zero-shot 67.8 169.4 62.8 47.8 492.8 399.9 56.0 332.5 232.1 DeepLabV3s Zero-shot 56.4 287.3 166.5 60.8 156.2 55.2 58.4 276.5 183.3 Multi-class Zero-shot 48.4 490.3 371.2 47.8 492.8 399.9 53.8 410.4 326.1 Swinunetr Zero-shot 53.3 231.3 105.8 49.6 179.1 63.4 55.6 226.3 110.2 Segmenter Zero-shot 49.4 251.1 141.1 49.4 261.1 117.7 55.3 283.3 180.5 GLAM Zero-shot(VM) 67.8 172.6 61.3 69.5 190.6 59.0 62.5 229.3 124.9 GLAM Zero-shot(VH) 69.3 160.7 52.3 69.8 194.0 64.1 62.9 226.7 118.7
Method Global sclerosis Haylinosis Mesangial lysis Microaneurysm Dice HD MSD Dice HD MSD Dice HD MSD Dice HD MSD U-Nets H2H 91.3 120.7 23.3 49.9 391.0 355.7 49.2 455.2 375.4 49.4 422.9 347.8 DeepLabV3s H2H 95.0 69.6 15.6 54.1 313.5 241.7 54.1 336.7 200.6 55.7 235.0 137.3 Multi-class H2H 92.6 103.8 17.7 49.9 391.0 355.7 49.2 521.6 441.8 49.4 422.9 347.8 Swinunetr H2H 90.4 131.1 26.1 49.9 314.8 285.3 49.2 382.1 273.6 49.5 336.7 252.5 Segmenter H2H 83.3 151.0 36.6 50.2 287.1 233.6 49.2 483.8 406.6 50.5 262.6 136.1 GLAM H2H 94.3 79.8 13.9 49.9 391.0 355.7 49.2 521.6 441.8 50.5 368.6 294.7 GLAM M&H2H 94.1 76.3 14.3 57.8 181.5 85.5 54.6 284.7 213.5 51.0 241.7 162.4
Method Nodular sclerosis Segmental sclerosis Average Dice HD MSD Dice HD MSD Dice HD MSD U-Nets H2H 67.0 158.6 55.8 47.8 492.8 399.9 59.1 340.2 259.7 DeepLabV3s H2H 70.9 151.4 58.9 52.1 331.0 220.6 63.7 239.5 145.8 Multi-class H2H 48.4 490.3 371.2 47.8 492.8 399.9 56.2 403.7 322.3 Swinunetr H2H 65.0 183.5 49.2 47.8 492.8 399.9 58.6 306.8 214.4 Segmenter H2H 49.2 310.3 202.0 45.0 283.0 177.5 54.6 296.3 198.7 GLAM H2H 76.5 152.2 38.8 68.5 181.2 57.0 64.8 282.4 200.3 GLAM M&H2H 76.8 145.9 37.0 70.1 200.9 63.5 67.3 188.5 96.0
@@ 4.2 Mouse & Human-to-Human Transfer Learning
Finally, we explored the model’s ability to generalize across species by training it on a combined dataset of both mouse and human samples. This hybrid training approach aims to leverage the strengths of both datasets to improve prediction accuracy on human samples. The performance of this model, detailed in Table 3, shows metrics for six lesion classes across the entire human dataset. According to our experimental results, the GLAM MH2H model not only successfully transfers learning across species but also outperforms most baseline human-to-human models in predicting different lesion classes. This highlights the potential benefits of integrating cross-species data to enhance model performance in complex prediction tasks.
@@ 5 Conclusion
In this study, we introduced a deep learning approach for fine-grained segmentation of human kidney lesions by adapting a mouse model, focusing on mouse-to-human transfer learning for glomerular lesions. For zero-shot transfer learning, we developed GLAM zero-shot models, designed to generate precise masks for lesion segmentation. Experimental results indicated that our zero-shot model surpassed baseline methods in segmenting glomerular lesions. For transfer learning with a hybrid dataset, we trained models using both mouse and human samples to predict on a human dataset. The results demonstrated that the hybrid model generally outperformed baseline methods and achieved satisfactory performance in predicting human datasets.
Acknowledgments and Disclosure of Funding
This research was supported by NIH R01DK135597(Huo), DoD HT9425-23-1-0003(HCY), NIH NIDDK DK56942(ABF). This work was also supported by Vanderbilt Seed Success Grant, Vanderbilt Discovery Grant, and VISE Seed Grant. This project was supported by The Leona M. and Harry B. Helmsley Charitable Trust grant G-1903-03793 and G-2103-05128. This research was also supported by NIH grants R01EB033385, R01DK132338, REB017230, R01MH125931, and NSF 2040462. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant.
@@ Appendix A Mouse-to-Mouse Supervised Learning
For evaluating the performance of prediction on mouse, under mouse-to-mouse (M2M) scenario, the model was trained using the mouse training dataset over 200 epochs and its performance was assessed on the mouse testing dataset. As shown in Table 4,the performance metrics for the segmentation of each glomerular lesion class over the entire mouse dataset. And Fig.5 presents the qualitative result about the performance of different methods on the multi-label mouse dataset. The experimental results show that GLAM M2M as a single multi-label model can achieve better performance on prediction classes of glomerular lesions than baseline methods. such as CNNs-based.DeepLabV3 and Transformers-based (i.e.Swin-unetr).
Figure 5: Qualitative Results of different methods on nine mice lesions for diabetic nephropathy. This figure displays the qualitative outcomes of various segmentation methods for nine classes of glomerular lesions. The first column features the original, unannotated images, while the second column shows the manual segmentation results. Subsequent columns are dedicated to the results from baseline models and the GLAM Mice-in-domain approach, respectively.
@@
Method Global sclerosis Haylinosis Mesangial lysis Microaneurysm Dice HD MSD Dice HD MSD Dice HD MSD Dice HD MSD U-Nets M2M 81.5 169.9 43.6 70.5 156.0 68.6 49.4 432.0 394.5 71.6 145.8 59.4 DeepLabV3s M2M 87.9 87.4 21.9 68.5 128.1 91.5 52.0 377.2 343.3 69.5 110.0 46.3 Multi-class M2M 87.6 123.0 30.5 49.9 391.9 368.5 49.4 432.0 394.5 49.0 443.9 371.6 Swinunetr M2M 84.3 128.1 33.5 73.3 90.6 45.1 50.8 164.4 88.8 70.5 136.9 36.1 Segmenter M2M 79.8 141.0 34.3 50.8 341.3 314.2 50.2 236.0 183.2 66.8 218.4 64.3 GLAM M2M 90.0 104.8 25.3 76.6 81.2 29.5 62.0 148.3 79.3 76.3 132.2 37.1
Method Nodular sclerosis Segmental sclerosis Average Dice HD MSD Dice HD MSD Dice HD MSD U-Nets M2M 62.3 205.4 135.0 47.4 475.0 372.0 63.8 264.0 178.9 DeepLabV3s M2M 70.1 126.7 62.9 66.0 143.9 54.8 69.0 162.2 103.5 Multi-class M2M 49.2 438.6 369.2 47.4 475.0 371.9 55.4 384.1 317.7 Swinunetr M2M 66.8 108.2 32.4 49.3 128.7 47.9 65.8 126.2 47.3 Segmenter M2M 56.5 166.0 59.0 51.5 202.2 80.1 59.3 217.5 122.5 GLAM M2M 77.7 108.3 37.2 78.2 145.9 39.6 76.8 120.1 41.3

### v2

# GLAM: Glomeruli Segmentation for Human Pathological Lesions using Adapted Mouse Model

@@ Abstract
keywords:
glomerulus, whole-slide image, glomerular lesion, segmentation, transfer learing
@@ 1 INTRODUCTION
Recent advancements such as  [1, 2, 3] in image processing and deep learning have significantly influenced the enhancement of computer-assisted analysis of medical images , particularly in the quantification of animal tissues via WSI. Within the domain of renal pathology, which is characterized by its unique challenges in pathology image analysis, the glomerulus is identified as critical functional units essential for clinical assessments, according to  [4, 5, 6].
Considerable research including [7, 8] has investigated the use of deep learning in the quantification of glomeruli within the field of renal pathology, employing technologies such as convolutional neural networks (CNNs) like [9, 10, 11, 12]. Interestingly, for glomerulus segmentation task, mouse-to-human transfer learning has been proven to be practical by [13, 6]. Nonetheless, these studies like [14] have devoted relatively little attention to the specific area of glomerular lesions. Glomerular lesions, which signify glomerular damage, serve as primary markers for various renal pathologies, according to [15]. Nonetheless, the segmentation of glomerular lesions, owing to their minuscule size and the variability encountered in intra- and inter-observer analyses, presents a more formidable challenge than the segmentation of healthy glomeruli. The analysis of glomerular lesions in humans is a fundamental aspect of nephrology, offering critical insights for diagnosis, understanding disease mechanisms, guiding treatment decisions, assessing prognosis, preventing kidney failure, and facilitating research into new treatments. Moreover, very few, if any, publicly accessible datasets are available for training segmentation models, which include a broad spectrum of lesion types in mouse kidney pathology images.
@@
Figure 1: Overview of the learning framework. This figure provides an overview of transfer learning for glomerular segmentation from mouse to human, highlighting the strenth of using mouse samples compared to human samples.
@@ 2 Method
We introduce the GLAM network, inspired from the research by [16], tailored for our glomerular lesion segmentation tasks. GLAM Zero-shot network is a unified segmentation network that leverages a residual U-Net architecture to segment various lesion classes within partially labeled images for pathology analysis. The comprehensive structure of the GLAM Zero-shot network is detailed in Figure referred to as Fig.2.
Unlike from fully labeled dataset, each image included in partially labeled dataset contains the annotations of only a specific class of glomerular lesion. We assign each segmentation on each class of lesion as a task for task-awareness for task-awareness, by encoding the task as a $m$-dimensional one-hot vector by [17]. The $m$ is the number of lesion classes, The encoding calculation for $T_{k}$, a class-aware vector of $i$th class of lesion is shown as follows:
@@
Dynamic filter generation was introduced by [18] to generate the kernels specialized to a particular class of lesion segment tasks. The image feature $F$ is aggregated by a global average pooling (GAP) with $T_{k}$. The kernel parameters $\omega$ are computed as follows:
@@
Table 1: Data Collection
Samples Size Scale #Patch
GS HN ML MA NS SS Sum
Mouse $512^{2}$ 80×80$\times$80 × 346 187 73 198 328 193 1325
Human $512^{2}$ 40×40$\times$40 × 370 36 23 57 229 72 787
@@ 3 Experiments
Figure 3: Lesion Classes. This figure shows fine-grained classes of mouse and human glomerular images that are stained with PAS. Glomeruli with lesions: global sclerosis(GS), Hyalinosis(HN), mesangial lysis(ML), microaneurysm(MA), nodular sclerosis(NS), and segmental sclerosis(SS).
@@ 3.1 Data description
The datasets utilized in this study, as detailed in Table.1, encompass both a mouse glomerular lesion dataset and a human glomerular lesion dataset. The mouse dataset collect patches from whole-slide images (WSIs) obtained from 17 mouse kidneys, spanning 6 lesion classes. The human dataset consists of patches derived from WSIs of 19 patients, also spanning 6 lesion classes. Initially, each patch in both datasets measured 1024x1024 pixels, but they were subsequently resized to 512x512 pixels and each patch contains at least one type of lesion. Based on the inherent size differences between human and mouse glomeruli. Human glomeruli are generally larger than those of mice, so to achieve a similar level of detail and ensure consistent visualization of lesions across species, we conducted our analysis at different magnifications: 40x for human samples and 80x for mouse samples, ensuring similar visualization and assessment of the lesions.
Method Global sclerosis Hyalinosis Mesangial lysis Microaneurysm Dice HD MSD Dice HD MSD Dice HD MSD Dice HD MSD U-Nets Zero-shot 70.6 185.0 51.6 49.9 352.4 317.0 49.2 521.6 441.8 50.9 273.5 119.7 DeepLabV3s Zero-shot 78.2 164.5 47.3 52.7 351.7 282.1 52.4 359.1 280.6 49.9 340.3 268.2 Multi-class Zero-shot 78.0 143.5 40.2 49.9 391.0 355.7 49.2 521.6 441.8 49.4 422.9 347.8 Swinunetr Zero-shot 81.5 147.6 37.3 50.1 319.0 251.5 49.3 225.1 94.2 49.5 255.6 109.1 Segmenter Zero-shot 82.5 146.6 36.6 50.4 267.8 224.7 49.2 521.6 441.8 50.7 251.6 121.2 GLAM Zero-shot(VM) 79.3 137.2 39.8 49.9 389.4 348.6 57.8 244.8 118.5 50.5 240.9 122.2 GLAM Zero-shot(VH) 82.0 125.3 35.9 49.9 387.5 351.9 56.2 248.9 96.8 50.2 243.8 111.4
Method Nodular sclerosis Segmental sclerosis Average
Dice HD MSD Dice HD MSD Dice HD MSD
U-Nets Zero-shot 67.8 169.4 62.8 47.8 492.8 399.9 56.0 332.5 232.1
DeepLabV3s Zero-shot 56.4 287.3 166.5 60.8 156.2 55.2 58.4 276.5 183.3
Multi-class Zero-shot 48.4 490.3 371.2 47.8 492.8 399.9 53.8 410.4 326.1
Swinunetr Zero-shot 53.3 231.3 105.8 49.6 179.1 63.4 55.6 226.3 110.2
Segmenter Zero-shot 49.4 251.1 141.1 49.4 261.1 117.7 55.3 283.3 180.5
GLAM Zero-shot(VM) 67.8 172.6 61.3 69.5 190.6 59.0 62.5 229.3 124.9
GLAM Zero-shot(VH) 69.3 160.7 52.3 69.8 194.0 64.1 62.9 226.7 118.7
@@ 3.2 Implementation details
We designed the image pool by incorporating strategies from Cycle-GAN by [19] for each category, with a batch size of four and an image pool size of eight. Excess images in the pool, when exceeding the batch size, were removed and input into the network.
@@
We compared the introduced network to baseline models, including: (1) multiple individual U-Net models (U-Nets) by  [20], (2) multiple individual DeepLabv3 models (DeepLabv3s) by  [21], and (3) a multi-class segmentation model for partially labeled datasets by  [22] for renal pathology quantification. Additionally, the performance of the network was evaluated against transformer baselines (4) a hybrid neural network architecture that combines the Swin Transformer with the U-Net transformer encoder for enhanced medical image segmentation (Swin-Unetr) by [23], and (5) an approach to semantic segmentation based on the Vision Transformer (Segmenter) by [24]. All of the parameter settings are followed by original paper.
Figure 4: Qualitative Results of different methods on human glomerular lesions. This figure displays the qualitative outcomes of various segmentation methods for six classes of glomerular lesions. The first column features the original, unannotated images, while the second column shows the manual segmentation results. Subsequent columns belong to three section: ”Mouse-to-human”,”Human-to-human” and ”Mouse&Human-to-human”, where VM refers to model selection utilizing validation results from mouse data, whereas VH indicates model selection based on validation results from human data. The bold mark indicates the best performance.
Method Global sclerosis Hyalinosis Mesangial lysis Microaneurysm Dice HD MSD Dice HD MSD Dice HD MSD Dice HD MSD U-Nets H2H 91.3 120.7 23.3 49.9 391.0 355.7 49.2 455.2 375.4 49.4 422.9 347.8 DeepLabV3s H2H 95.0 69.6 15.6 54.1 313.5 241.7 54.1 336.7 200.6 55.7 235.0 137.3 Multi-class H2H 92.6 103.8 17.7 49.9 391.0 355.7 49.2 521.6 441.8 49.4 422.9 347.8 Swinunetr H2H 90.4 131.1 26.1 49.9 314.8 285.3 49.2 382.1 273.6 49.5 336.7 252.5 Segmenter H2H 83.3 151.0 36.6 50.2 287.1 233.6 49.2 483.8 406.6 50.5 262.6 136.1 GLAM H2H 94.3 79.8 13.9 49.9 391.0 355.7 49.2 521.6 441.8 50.5 368.6 294.7 GLAM M&H2H 94.1 76.3 14.3 57.8 181.5 85.5 54.6 284.7 213.5 51.0 241.7 162.4
Method Nodular sclerosis Segmental sclerosis Average
Dice HD MSD Dice HD MSD Dice HD MSD
U-Nets H2H 67.0 158.6 55.8 47.8 492.8 399.9 59.1 340.2 259.7
DeepLabV3s H2H 70.9 151.4 58.9 52.1 331.0 220.6 63.7 239.5 145.8
Multi-class H2H 48.4 490.3 371.2 47.8 492.8 399.9 56.2 403.7 322.3
Swinunetr H2H 65.0 183.5 49.2 47.8 492.8 399.9 58.6 306.8 214.4
Segmenter H2H 49.2 310.3 202.0 45.0 283.0 177.5 54.6 296.3 198.7
GLAM H2H 76.5 152.2 38.8 68.5 181.2 57.0 64.8 282.4 200.3
GLAM M&H2H 76.8 145.9 37.0 70.1 200.9 63.5 67.3 188.5 96.0
@@ 4.2 Mouse & Human-to-Human Transfer Learning
Finally, we explored the model’s ability to generalize across species by training it on a combined dataset of both mouse and human samples. This hybrid training approach aims to leverage the strengths of both datasets to improve prediction accuracy on human samples. The performance of this model, detailed in Table 3, shows metrics for six lesion classes across the entire human dataset.
@@ 5 Conclusion
In this study, we introduced a deep learning approach for fine-grained segmentation of human kidney lesions by adapting a mouse model, focusing on mouse-to-human transfer learning for glomerular lesions. For zero-shot transfer learning, we developed GLAM zero-shot models, designed to generate precise masks for lesion segmentation. Experimental results indicated that our zero-shot model surpassed baseline methods in segmenting glomerular lesions. For transfer learning with a hybrid dataset, we trained models using both mouse and human samples to predict on a human dataset in order to leverages mouse samples as supplementary support to enhance performance in predicting human lesions. The results demonstrated that the hybrid model generally outperformed baseline methods and achieved satisfactory performance in predicting human datasets.
@@ Appendix A Mouse-to-Mouse Supervised Learning
For evaluating the performance of prediction on mouse, under mouse-to-mouse (M2M) scenario, the model was trained using the mouse training dataset over 200 epochs and its performance was assessed on the mouse testing dataset. As shown in Table 4,the performance metrics for the segmentation of each glomerular lesion class over the entire mouse dataset.
@@
Method Global sclerosis Hyalinosis Mesangial lysis Microaneurysm Dice HD MSD Dice HD MSD Dice HD MSD Dice HD MSD U-Nets M2M 81.5 169.9 43.6 70.5 156.0 68.6 49.4 432.0 394.5 71.6 145.8 59.4 DeepLabV3s M2M 87.9 87.4 21.9 68.5 128.1 91.5 52.0 377.2 343.3 69.5 110.0 46.3 Multi-class M2M 87.6 123.0 30.5 49.9 391.9 368.5 49.4 432.0 394.5 49.0 443.9 371.6 Swinunetr M2M 84.3 128.1 33.5 73.3 90.6 45.1 50.8 164.4 88.8 70.5 136.9 36.1 Segmenter M2M 79.8 141.0 34.3 50.8 341.3 314.2 50.2 236.0 183.2 66.8 218.4 64.3 GLAM M2M 90.0 104.8 25.3 76.6 81.2 29.5 62.0 148.3 79.3 76.3 132.2 37.1
Method Nodular sclerosis Segmental sclerosis Average
Dice HD MSD Dice HD MSD Dice HD MSD
U-Nets M2M 62.3 205.4 135.0 47.4 475.0 372.0 63.8 264.0 178.9
DeepLabV3s M2M 70.1 126.7 62.9 66.0 143.9 54.8 69.0 162.2 103.5
Multi-class M2M 49.2 438.6 369.2 47.4 475.0 371.9 55.4 384.1 317.7
Swinunetr M2M 66.8 108.2 32.4 49.3 128.7 47.9 65.8 126.2 47.3
Segmenter M2M 56.5 166.0 59.0 51.5 202.2 80.1 59.3 217.5 122.5
GLAM M2M 77.7 108.3 37.2 78.2 145.9 39.6 76.8 120.1 41.3
Figure 5: Qualitative Results of different methods on six mice lesions for diabetic nephropathy. This figure displays the qualitative outcomes of various segmentation methods for six classes of glomerular lesions. The first column features the original, unannotated images, while the second column shows the manual segmentation results. Subsequent columns are dedicated to the results from baseline models and the GLAM Mouse-to-Mouse approach, respectively.
Fig.5 presents the qualitative result about the performance of different methods on the mouse lesion dataset. The experimental results show that GLAM M2M as a single multi-label model can achieve better performance on prediction classes of glomerular lesions than baseline methods. such as CNNs-based.DeepLabV3 and Transformers-based (i.e.Swin-unetr).
Acknowledgements.
This research was supported by NIH R01DK135597(Huo), DoD HT9425-23-1-0003(HCY), NIH NIDDK DK56942
(ABF). This work was also supported by Vanderbilt Seed Success Grant, Vanderbilt Discovery Grant, and VISE Seed Grant. This project was supported by The Leona M. and Harry B. Helmsley Charitable Trust grant G-1903-03793 and G-2103-05128. This research was also supported by NIH grants R01EB033385, R01DK132338, REB017230, R01MH125931, and NSF 2040462. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant. This works was also supported by NSF NAIRR Pilot Award NAIRR240055.